# Remarks on the inverse Littlewood conjecture

### v1

@@ Abstract.
The Littlewood conjecture, proven by Konyagin and McGehee-Pigno-Smith in the 1980s, states that if $A\subset\mathbb{Z}$ is a finite set of integers with $\lvert A\rvert=N$ then $\|\widehat{1_{A}}\|_{1}\geqslant c\log N$ for some absolute constant $c>0$. We explore what structure $A$ must have if $\|\widehat{1_{A}}\|_{1}\leqslant K\log N$ for some constant $K$. Under such an assumption we prove, for instance, that $A$ contains a subset $A^{\prime}\subseteq A$ with $\lvert A\rvert\geqslant N^{0.99}$ such that $\lvert A^{\prime}+A^{\prime}\rvert\ll K^{O(1)}\lvert A^{\prime}\rvert$. As a consequence, for any $k\geqslant 3$, if $N$ is sufficiently large depending on $k$ and $K$, then $A$ must contain an arithmetic progression of length $k$. A byproduct of our analysis is a (slightly) improved bound for the constant $c$.
@@ 1.2. Bounds for the Littlewood conjecture
where $c=0.1422917\cdots$.
For comparison note that $4/\pi^{2}\approx 0.405$ so we are still about a factor of $3$ away from the optimal value. The value of $c$ is found as the maximum of a complicated two-variable expression and presumably has no nicer form.
@@
Note that if $A=\mathrm{supp}(f)$ and $B=\mathrm{supp}(g)$ then $\mathrm{supp}(f\ast g)\subseteq A+B$. All of the functions we consider will have finite support, so there are no issues of convergence. We note here the useful property that $\widehat{f\ast g}=\widehat{f}\cdot\widehat{g}$. It is convenient to use inner product notation, so that if $f,g:\mathbf{Z}\to\mathbf{C}$ we write
@@ Acknowledgements
TB is supported by a Royal Society University Research Fellowship. BG is supported by a Simons Investigator award. We thank Benjamin Bedert for many insightful and productive discussions concerning the inverse Littlewood problem.
@@ 2. McGehee-Pigno-Smith test functions
We then set $\mathcal{R}=\mathcal{R}_{J}$. The fact that $\mathcal{R}$ has the form (2.1) is a trivial induction.
@@
$\widehat{F_{ij}}:=\widehat{\mathcal{R}_{g_{j}}}\prod_{i<\ell\leqslant j}(\widehat{\mathcal{R}_{g_{\ell}}}+1),$
@@
The bound (2.3) now follows using Lemma 2.1 (2) and (3) and Parseval’s identity $\|\widehat{F_{ij}}\|_{2}=\|F_{ij}\|_{2}$. The fact that $\mathcal{R}$ is supported on $\mathbf{Z}_{\leqslant 0}$ follows from Lemma 2.1 (1). ∎
@@
Let $b>0$ be a real parameter, and let $J\in\mathbf{N}$. Let $A\subset\mathbf{Z}$ be a finite set and suppose that $A_{1},\dots,A_{J}$ are initial segments of $A$ with $A_{1}\subseteq\cdots\subseteq A_{J}$. Let $\lambda>1$ be a parameter, and suppose that $\left\lvert A_{j+1}\right\rvert\geqslant\lambda\left\lvert A_{j}\right\rvert$ for $j=1,\dots,J$. Then
$\|\widehat{1_{A}}\|_{1}\geqslant(1-e^{-b})\left(J-\frac{b\sqrt{2}}{\sqrt{\lambda}-1}\sum_{i=1}^{J}\omega[A_{i}]^{1/2}\right),$
where $\omega[\cdot]$ denotes normalised additive energy.
@@
By the Cauchy-Schwarz inequality, we thus have
$\big|\langle 1_{A},\mu_{i}\ast F_{ij}\rangle\big|=\big|\langle 1_{A_{i}}\ast\mu_{i}^{\circ},F_{ij}\rangle\big|\leqslant 2^{1/2}b|A_{j}|^{-1}\|1_{A_{i}}\ast\mu_{i}^{\circ}\|_{2}.$ (3.3)
@@
$\|1_{A_{i}}\ast\mu_{i}^{\circ}\|_{2}=\frac{E(A_{i})^{1/2}}{|A_{i}|}.$
@@
$\Big|\sum_{j=1}^{J}\langle D_{j},1_{A}\rangle\Big|\leqslant c2^{1/2}b\sum_{1\leqslant i<j\leqslant J}\frac{E(A_{i})^{1/2}}{|A_{i}||A_{j}|^{1/2}}.$
@@
$\|\widehat{1_{A}}\|_{1}\geqslant(1-e^{-b})J\Big(J-2^{1/2}b\sum_{i=1}^{J}\frac{E(A_{i})^{1/2}}{\left\lvert A_{i}\right\rvert}\sum_{i<j\leqslant J}\left\lvert A_{j}\right\rvert^{-1/2}\Big).$
By the assumption in the lemma we have $|A_{j}|^{-1/2}\leqslant\lambda^{(i-j)/2}|A_{i}|^{-1/2}$ for $j\geqslant i$. Summing the geometric series yields
@@ 4. Improved constants for Littlewood’s conjecture
$E(A,B)=\#\{(a_{1},b_{1},a_{2},b_{2})\in A\times B\times A\times B\}\leqslant nm^{2}\big(1-\frac{m}{3n}\big)+\frac{m}{3}.$
@@
$r(a_{i}+b_{j})\leqslant 1+\min(m-i,j-1)+\min(n-j,i-1).$ (4.1)
Indeed, the constant $1$ counts the representation $a_{i}+b_{j}$ itself. The $\min(m-i,j-1)$ term is an upper bound for representations $a_{i}+b_{j}=a_{i^{\prime}}+b_{j^{\prime}}$ with $i<i^{\prime}\leqslant m$, which must have also have $1\leqslant j^{\prime}<j$; for each $i^{\prime}$ (or $j^{\prime}$) there is at most one such representation. The $\min(n-j,i-1)$ term is an upper bound for representations $a_{i}+b_{j}=a_{i^{\prime}}+b_{j^{\prime}}$ with $1\leqslant i^{\prime}<i$ (and hence $j<j^{\prime}\leqslant n$). Now sum (4.1) over $1\leqslant i\leqslant m$ and $1\leqslant j\leqslant n$. The left-hand side becomes $E(A,B)$, and (after a computation) the right-hand side becomes $nm^{2}\big(1-\frac{m}{3n}\big)+\frac{m}{3}$. ∎
@@
Let $A_{1}$ be the set consisting of the smallest $\lfloor\log N\rfloor$ elements of $A$. In general, if $A_{1},\ldots,A_{j}$ have been constructed, let $A_{j+1}$ be the set of the smallest $\lceil\lambda\left\lvert A_{j}\right\rvert\rceil$ elements of $A$ and continue. Otherwise, we set $J=j$ and halt. Furthermore we have $\left\lvert A_{j+1}\right\rvert\geqslant\lambda\left\lvert A_{j}\right\rvert$ for $j=1,\dots,J$. Additionally, $\left\lvert A_{j+1}\right\rvert\leqslant\lambda\left\lvert A_{j}\right\rvert+1$ for $j=1,\dots,J$, and hence by induction
@@
$f(b,\lambda)=\frac{1-e^{-b}}{\log\lambda}\left(1-\frac{2b}{\sqrt{3}(\sqrt{\lambda}-1)}\right).$
@@
$b=1.3766505\cdots\quad\textrm{and}\quad\lambda=36.1127893\cdots.\qed$
@@ 5. A structural inverse result
This follows immediately from Proposition 3.1 with $b=1$ and $1/\lambda=1-\eta$, so that $\frac{1}{\sqrt{\lambda}-1}\ll\eta^{-1}$. ∎
@@
We construct nested initial segments $A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{J}$ of $A$ in the following fashion. Let $A_{1}$ be the smallest $\lfloor N^{1-\delta}\rfloor$ elements of $A$. In general, if $A_{1},\ldots,A_{j}$ have been constructed, then let $A_{j+1}$ be the set of the smallest $\lceil(1-\eta)^{-1}\left\lvert A_{j}\right\rvert\rceil$ elements of $A$ if $A$ has this many elements, or else halt with $j=J$. By construction it is clear that $\left\lvert A_{j}\right\rvert\leqslant(1-\eta)\left\lvert A_{j+1}\right\rvert$ for all $j\in\{1,\dots,J\}$. Furthermore, $\left\lvert A_{j+1}\right\rvert\leqslant(1-\eta)^{-1}\left\lvert A_{j}\right\rvert+1$ for $j\in\{1,\dots,J\}$, and hence by the same induction leading to (4.2) we have

### v2

@@ Abstract.
The Littlewood conjecture, proven by Konyagin and McGehee-Pigno-Smith in the 1980s, states that if $A\subset\mathbb{Z}$ is a finite set of integers with $\lvert A\rvert=N$ then $\|\widehat{1_{A}}\|_{1}\geqslant c\log N$ for some absolute constant $c>0$. We explore what structure $A$ must have if $\|\widehat{1_{A}}\|_{1}\leqslant K\log N$ for some constant $K$. Under such an assumption we prove, for instance, that $A$ contains a subset $A^{\prime}\subseteq A$ with $\lvert A^{\prime}\rvert\geqslant N^{0.99}$ such that $|A^{\prime}+A^{\prime}|\ll K^{O(1)}|A^{\prime}|$. As a consequence, for any $k\geqslant 3$, if $N$ is sufficiently large depending on $k$ and $K$, then $A$ must contain an arithmetic progression of length $k$. A byproduct of our analysis is a (slightly) improved bound for the constant $c$.
@@ 1.2. Bounds for the Littlewood conjecture
where $c=0.170934\cdots$.
For comparison note that $4/\pi^{2}\approx 0.405$ so we are still about a factor of $2.4$ away from the conjectured optimal value. The value of $c$ is found as the maximum of a complicated two-variable expression and presumably has no nicer form.
@@
Note that if $A=\mathrm{supp}(f)$ and $B=\mathrm{supp}(g)$ then $\mathrm{supp}(f\ast g)\subseteq A+B$. We note here the useful property that $\widehat{f\ast g}=\widehat{f}\cdot\widehat{g}$. It is convenient to use inner product notation, so that if $f,g:\mathbf{Z}\to\mathbf{C}$ we write
@@ Acknowledgements
TB is supported by a Royal Society University Research Fellowship. BG is supported by a Simons Investigator award. We thank Benjamin Bedert for many insightful and productive discussions concerning the inverse Littlewood problem, and the referee for observing that a minor modification to the argument gives a larger value of $c$ than we had originally claimed.
@@ 2. McGehee-Pigno-Smith test functions
(Here $\mathcal{R}_{g_{i+1}}$ is the test function given by Lemma 2.1 with $f=g_{i+1}$.) We then set $\mathcal{R}=\mathcal{R}_{J}$. The fact that $\mathcal{R}$ has the form (2.1) is a trivial induction.
@@
$\widehat{F_{ij}}:=\widehat{\mathcal{R}_{g_{j}}}\prod_{i<\ell<j}(\widehat{\mathcal{R}_{g_{\ell}}}+1),$
@@
The bound (2.3) now follows using Lemma 2.1 (2) and (3) and Parseval’s identity $\|\widehat{F_{ij}}\|_{2}=\|F_{ij}\|_{2}$. To see that $F_{ij}$ is supported on $\mathbf{Z}_{\leqslant 0}$, observe that $F_{ij}$ is a sum of convolutions of functions $\mathcal{R}_{g_{\ell}}$, and each such function is supported on $\mathbf{Z}_{\leqslant 0}$ by Lemma 2.1 (1). ∎
@@
Let $b>0$ be a real parameter, and let $J\in\mathbf{N}$. Let $A\subset\mathbf{Z}$ be a finite set and suppose that $A_{1},\dots,A_{J}$ are initial segments of $A$ with $A_{1}\subseteq\cdots\subseteq A_{J}$. Let $\lambda>1$ be a parameter, and suppose that $\left\lvert A_{j+1}\right\rvert\geqslant\lambda\left\lvert A_{j}\right\rvert$ for $j=1,\dots,J-1$. Then
$\|\widehat{1_{A}}\|_{1}\geqslant(1-e^{-b})\left(J-\frac{b}{\sqrt{\lambda}-1}\sum_{i=1}^{J}(\omega[A_{i}]+\left\lvert A_{i}\right\rvert^{-1})^{1/2}\right),$
where $\omega[\cdot]$ denotes the normalised additive energy.
@@
By the Cauchy-Schwarz inequality, we thus have, recalling that $F_{ij}$ is supported on $\mathbf{Z}_{\leqslant 0}$,
$\big|\langle 1_{A},\mu_{i}\ast F_{ij}\rangle\big|=\big|\langle 1_{A_{i}}\ast\mu_{i}^{\circ},F_{ij}\rangle\big|\leqslant 2^{1/2}b|A_{j}|^{-1/2}\big\|(1_{A_{i}}\ast\mu_{i}^{\circ})1_{\mathbf{Z}_{\leqslant 0}}\big\|_{2}.$ (3.3)
@@
$\|(1_{A_{i}}\ast\mu_{i}^{\circ})1_{\mathbf{Z}_{\leqslant 0}}\|_{2}=\left\lvert A_{i}\right\rvert^{-1}\left(\frac{E(A_{i})+\left\lvert A_{i}\right\rvert^{2}}{2}\right)^{1/2}.$
@@
$\Big|\sum_{j=1}^{J}\langle D_{j},1_{A}\rangle\Big|\leqslant cb\sum_{1\leqslant i<j\leqslant J}\frac{(E(A_{i})+\left\lvert A_{i}\right\rvert^{2})^{1/2}}{|A_{i}||A_{j}|^{1/2}}.$
@@
$\|\widehat{1_{A}}\|_{1}\geqslant(1-e^{-b})\Big(J-b\sum_{i=1}^{J}\frac{(E(A_{i})+\left\lvert A_{i}\right\rvert^{2})^{1/2}}{\left\lvert A_{i}\right\rvert}\sum_{i<j\leqslant J}\left\lvert A_{j}\right\rvert^{-1/2}\Big).$
By the assumption in the proposition we have $|A_{j}|^{-1/2}\leqslant\lambda^{(i-j)/2}|A_{i}|^{-1/2}$ for $j\geqslant i$. Summing the geometric series yields
@@ 4. Improved constants for Littlewood’s conjecture
$E(A,B)=\#\{(a_{1},b_{1},a_{2},b_{2})\in A\times B\times A\times B:a_{1}+b_{1}=a_{2}+b_{2}\}\leqslant nm^{2}\big(1-\frac{m}{3n}\big)+\frac{m}{3}.$
@@
$r(a_{i}+b_{j})\leqslant 1+\min(n-i,j-1)+\min(m-j,i-1).$ (4.1)
Indeed, the constant $1$ counts the representation $a_{i}+b_{j}$ itself. The $\min(n-i,j-1)$ term is an upper bound for representations $a_{i}+b_{j}=a_{i^{\prime}}+b_{j^{\prime}}$ with $i<i^{\prime}\leqslant n$, which must have also have $1\leqslant j^{\prime}<j$; for each $i^{\prime}$ (or $j^{\prime}$) there is at most one such representation. The $\min(m-j,i-1)$ term is an upper bound for representations $a_{i}+b_{j}=a_{i^{\prime}}+b_{j^{\prime}}$ with $1\leqslant i^{\prime}<i$ (and hence $j<j^{\prime}\leqslant m$). Now sum (4.1) over $1\leqslant i\leqslant n$ and $1\leqslant j\leqslant m$. The left-hand side becomes $E(A,B)$, and (after a computation) the right-hand side becomes $nm^{2}\big(1-\frac{m}{3n}\big)+\frac{m}{3}$. ∎
@@
Let $A_{1}$ be the set consisting of the smallest $\lfloor\log N\rfloor$ elements of $A$. In general, if $A_{1},\ldots,A_{j}$ have been constructed, let $A_{j+1}$ be the set of the smallest $\lceil\lambda\left\lvert A_{j}\right\rvert\rceil$ elements of $A$ and continue. Otherwise, we set $J=j$ and halt. Furthermore we have $\left\lvert A_{j+1}\right\rvert\geqslant\lambda\left\lvert A_{j}\right\rvert$ for $j=1,\dots,J-1$. Additionally, $\left\lvert A_{j+1}\right\rvert\leqslant\lambda\left\lvert A_{j}\right\rvert+1$ for $j=1,\dots,J-1$, and hence by induction
@@
$f(b,\lambda)=\frac{1-e^{-b}}{\log\lambda}\left(1-(2/3)^{1/2}\frac{b}{\sqrt{\lambda}-1}\right).$
@@
$b=0.932199\cdots\quad\textrm{and}\quad\lambda=9.112038\cdots.\qed$
@@ 5. A structural inverse result
Noting that $\omega[A]\geqslant|A|^{-1}$ for any nonempty set $A$ of integers, this follows immediately from Proposition 3.1 with $b=1$ and $1/\lambda=1-\eta$, so that $\frac{1}{\sqrt{\lambda}-1}\ll\eta^{-1}$. ∎
@@
We construct nested initial segments $A_{1}\subseteq A_{2}\subseteq\cdots\subseteq A_{J}$ of $A$ in the following fashion. Let $A_{1}$ be the smallest $\lfloor N^{1-\delta}\rfloor$ elements of $A$. In general, if $A_{1},\ldots,A_{j}$ have been constructed, then let $A_{j+1}$ be the set of the smallest $\lceil(1-\eta)^{-1}\left\lvert A_{j}\right\rvert\rceil$ elements of $A$ if $A$ has this many elements, or else halt with $j=J$. By construction it is clear that $\left\lvert A_{j}\right\rvert\leqslant(1-\eta)\left\lvert A_{j+1}\right\rvert$ for all $j\in\{1,\dots,J-1\}$. Furthermore, $\left\lvert A_{j+1}\right\rvert\leqslant(1-\eta)^{-1}\left\lvert A_{j}\right\rvert+1$ for $j\in\{1,\dots,J-1\}$, and hence by the same induction leading to (4.2) we have